OpenStudy (tabdo1226):

In his first five basketball games, Neal scored 10, 11, 11, 12, and 16 points. What is the mean absolute deviation for this set of scores? points

OpenStudy (tabdo1226):

@Bee_see

OpenStudy (tabdo1226):

need help more medals

OpenStudy (tabdo1226):

will fan and medal

OpenStudy (tabdo1226):

ignore the points because it's the answer

OpenStudy (tabdo1226):

not the answer but like 12 points

OpenStudy (tabdo1226):

is it 14.4

OpenStudy (amorfide):

find the mean by adding them all up and dividing by the amount of numbers that are there so if I have 5,6,7,8 I would add them up and divide by 4 now we subtract the mean from each of the numbers and take the absolute value add them together then divide by how many there are so 5,6,7,8 5+6+7+8=26 26/4=6.5 now I do 5-6.5 6-6.5 7-6.5 8-6.5 taking the absolute value of each one means you will take the positive version of that answer so -1 becomes 1 |-1|=1 then you add them all up so we get 1.5+0.5+0.5+1.5=4 then we divide by how many there are which is 4 4/4=1 the equation I have explained it \[\frac{ \sum_{}^{}|x=\mu | }{ N }\] where µ is the mean of the given numbers and N is the amount there are

OpenStudy (tabdo1226):

@Bee_see

OpenStudy (tabdo1226):

so it's 14.4

OpenStudy (tabdo1226):

right @amorfide

OpenStudy (tabdo1226):

is it 14.4

OpenStudy (tabdo1226):

http://inquus.org/updates/536c755ae4b03726c6600f6d

OpenStudy (tabdo1226):

@amorfide is it 14.4

OpenStudy (amorfide):

8/5 = 1.6

OpenStudy (tabdo1226):

so 1.6

OpenStudy (amorfide):

work it out

OpenStudy (tabdo1226):

@amorfide is it 1.6

OpenStudy (amorfide):

yes

OpenStudy (tabdo1226):

thank you i'll subit my test and tell you my score

OpenStudy (tabdo1226):

*sumbit

OpenStudy (tabdo1226):

90% thank you

OpenStudy (tabdo1226):

medal and fan